# Large Language Model Lateral Spear Phishing: A Comparative Study in Large-Scale Organizational Settings

### v1

@@ 5.4 RQ4: Effectiveness of Phishing Training and Warnings
Email Template Email Recipients Emails Opened Link Clicked Data Entered
Targeted Internal Information 208 162 (77.88%) 75 (36.05%) 51 (24.51%)
Group Internal Information 222 181 (81.53%) 89 (40.09%) 46 (20.72%)
External Information 216 85 (39.35%) 0 (0.00%) 0 (0.00%)
Total 646 428 (66.25%) 164 (25.38%) 97 (15.01%)
TABLE V: LLM crafted lateral phishing emails on the subset of users who entered data in the human phishing email experiment and then received phishing training. Percentage shown is percent relative to the number of email recipients.
5.5 RQ5: Identifying and Building Defense against LLM-Generated Phishing Emails
To answer RQ5, we evaluate some machine-generated text detection models on emails written by an LLM. To test the capability of existing machine-generated text detection systems to detect LLM generated emails, we first utilize a dataset of 5,000 human written email samples from the dataset by Chakraborty (2023) [32]. Each email is then rewritten with the Vicuna 13B v1.5 model from HuggingFace [33]. The Vicuna model was introduced by Zheng et al. [34] and a key feature of Vicuna is that it utilizes instruction tuning, which refines the model’s ability to understand and adhere to specific directives [35]. For the Vicuna model implementation we use 4-bit quantization and a context window of 16,384 tokens. In our experiment, Vicuna is consistently prompted to revise each email sample, resulting in pairs of human and machine written texts. The prompt used for rewriting the human written emails is as follows: ”Revise the following email for me. Only reply with the revised email and do not print anything else. Also, Do not include a subject in your answer. Here is the email text: [EMAIL TEXT]”.
We test four models from the work of Pu et al. [36]: BERT-Defense, GLTR-BERT, GLTR-GPT2, and RoBERTa-Defense. GLTR-BERT and GLTR-GPT2 distinguish synthetic from genuine text by leveraging token probability distributions. These methods are based on the premise that synthetic text often incorporates tokens with high probability rankings in language models. GLTR derives features from token rankings in the Top-10, Top-100, and Top-1000, which is then processed by a logistic regression classifier. GLTR-BERT employs BERT as its base language model, while GLTR-GPT2 uses GPT2-XL. Their training involved 4,000 articles from WebText [37] for the human class and GPT2-XL generations for the synthetic class. BERT-Defense is a binary classifier built upon the BERT-Large pretrained language model through the addition of a binary classification layer. Inspired by BERT-Defense, RoBERTa-Defense uses the RoBERTa language model as a base with a layer for binary classification. BERT-Defense was trained on 10,000 articles each from WebText for real articles and GPT2-Large for synthetic ones, while RoBERTa-Defense used the RealNews dataset and GROVER [38] generations, with 5,000 articles from each for genuine and synthetic texts, respectively.
To show that LLM generated emails can be detected, we first evaluate the four detectors using their base trained weights to test their ability to generalize to the phishing email domain without any modifications. Then, we perform a fine-tuning experiment following the work of Pu et al. [36] on the models to show that the model’s performance can be significantly improved when exposed to data in the email domain. We fine-tune BERT-Defense and RoBERTa-Defense using samples from the target distribution using 100 samples, 50 from both the human and machine class. The evaluation metrics we present are the precision, recall, and F1 score with respect to the machine class. The first 4 rows of Table VI show varied performance of the base models, with RoBERTa-Defense achieving the highest F1 score of 67.1, and GLTR-GPT2 performing poorly with an F1 score of 6.6, suggesting challenges of the base models to generalize to LLM generated emails. In rows 5-8 of Table VI, we present the results of our fine-tuning experiment and note that all models exhibit a substantial improvement in performance. BERT-Defense and GLTR-GPT2, which had the worst base model performance showed the largest improvements in this fine-tuning experiment, with both models having an increase in F1 score by more than 69%.
We also build a model, T5-LLM Email Defense, using the T5 (Text-to-Text Transfer Transformer) encoder [39] to distinguish between human and LLM generated email texts. The model architecture integrates the T5 encoder model with an additional classification layer to perform binary classification. We choose to use the T5 encoder model over the full T5 model since binary classification requires understanding and categorizing of input text, which is only a function of the encoder. T5-LLM Email Defense is trained on our dataset of human and LLM generated emails, using the AdamW optimizer and a linear learning rate scheduler. The performance results are in the bottom row of Table VI, showing the highest F1 score of 98.96.
Model Data Source P R F1
BERT-Defense WebText/GPT2-Large 33.3 15.0 20.7
RoBERTa-Defense RealNews/GROVER 66.0 68.2 67.1
GLTR-BERT WebText/GPT2-XL 100 22.4 36.6
GLTR-GPT2 WebText/GPT2-XL 94.4 3.4 6.6
BERT-DefenseEmail LLM Generated Emails 87.6 92.0 89.8
RoBERTa-DefenseEmail 97.1 100 98.5
GLTR-BERTEmail 86.5 90.0 88.2
GLTR-GPT2Email 80.0 72.0 75.8
T5-LLMEmail Defense (ours) LLM Generated Emails 99.03 98.89 98.96
TABLE VI: Performance metrics of existing base detectors, fine-tuned versions of the existing base detectors, and our own detector with respect to the machine class.
5.6 Overall discussion and findings
@@
The after action review conducted by the cyber operations team yielded crucial insights into the efficacy of current email filtering platforms and the level of employee preparedness in identifying phishing emails, particularly those generated by LLMs. Given the sophistication of LLM-generated emails, which often lack grammatical errors and appear highly credible, it’s imperative to enhance the security awareness training program. This updated training should focus on educating employees to discern subtle cues, such as the overall context of the email and the nature of the request. Alongside recognizing these nuances, it is equally important for employees to be alert to common phishing red flags like the urgency of requests and the origin of URLs or attachments.
In parallel with employee training, it is critical to refine the email filtering platforms to detect phishing attempts more effectively. This includes going beyond traditional red flags and incorporating advanced detection techniques such as mail header analysis, monitoring unusual sender behaviors, comprehensive URL and attachment analysis, content scrutiny, recipient analysis, sender reputation assessment, and identifying spoofed display names or domains. A crucial addition to these methods is the implementation of LLM detectors that can identify inconsistencies in writing style and deviations from the sender’s typical communication pattern or discuss unfamiliar topics.
When an LLM-generated email is detected, proactive measures should be taken to mitigate its potential impact. One potentially effective strategy is to automatically tag such emails, informing the recipient of the potential risks. This tag would act as an immediate visual cue, alerting the recipient that the email has been flagged as being potentially LLM-generated, thus warranting extra scrutiny. Email tagging has been investigated in previous research as it augments users ability to more quickly categorize content [44, 45]. Currently within the organization, a similar tagging methodology is used to indicate emails that come from outside the organization, assigning an [EXTERNAL] tag to all such emails. Such tags would not only serve as a direct warning but also reinforce the training employees receive, bridging the gap between theoretical knowledge and practical application. This dual approach of enhancing both technological defenses and employee vigilance could be part of a defense strategy against the evolving landscape of LLM-generated phishing threats. Future research could investigate email tagging for LLM-generated content to see if these tags could heighten user awareness to LLM phishing threats.
6 Related Prior Work
Recently, concerns around LLMs being used for adversarial purposes, particularly in the context of phishing attacks, have risen significantly. A recent report by Google Cloud underscores that LLMs can be used to create highly legitimate-seeming content, including voice and video, making it more challenging to identify misspellings, grammar errors, and cultural inaccuracies in phishing emails and messages [57]. LLMs can also effectively translate and refine content, further complicating the detection of phishing attempts based on language use. The report points out that with generative AI, attackers can scale their operations and target a wide range of individuals with personalized and convincing emails, using data like names, organizations, job titles, and even health information. The work of Hazell et al. studied the effectiveness of LLMs toward creating realistic and cost-effective spear phishing emails to over 600 British Members of Parliament [58]. Their work also notes the possibility of bypassing safeguards in LLMs through basic prompt engineering, and emphasize the need for robust governance measures to prevent misuse. Another recent study [59] involved involved two groups of participants, exposed to either human-crafted or GPT-3 crafted phishing emails. Participants were asked to determine the authenticity of the emails, with feedback provided in the second round. The results indicated that human-crafted emails were more effective in deceiving people than those created by GPT-3, even after participants received training across various cognitive biases. Interestingly, participants felt more confident identifying phishing emails when they were human-crafted compared to those generated by GPT-3.
7 Conclusion
This study represents a significant advancement in understanding the evolving landscape of cyber threats, particularly in the context of Large Language Models (LLMs) and their role in spear phishing attacks. Through a comprehensive 11-month investigation involving roughly 9000 employees at a large public university, we have uncovered critical insights into the efficacy and dangers posed by LLM-generated lateral spear phishing emails. Our findings highlight several key areas of concern and opportunities for improving cybersecurity measures in large organizations.
Our research revealed that a notable percentage of the university’s staff was vulnerable to these AI-crafted attacks, with around 10% of email recipients entering their login credentials in response to these sophisticated phishing attempts. This high success rate of LLM-generated emails underscores the need for heightened awareness and training regarding these new forms of cyber threats. Secondly, our study exposes the limitations of current anti-phishing infrastructures in detecting and preventing LLM-generated phishing attacks. We implemented a machine learning-based detection technique specifically tailored to identify LLM-generated phishing emails. With an F1-score of 98.96, this technique was shown to be effective, paving the way for its integration into existing cybersecurity frameworks. Finally, the insights gained from the university’s cyber operations team are invaluable. They highlight the critical need for updated training and educational approaches focusing on LLM phishing threats. The evolving nature of cyber threats demands a dynamic and responsive approach to cybersecurity education and training, ensuring that organizations and their employees are well-equipped to recognize and counter these sophisticated attacks.

### v2

@@ VI-B RQ2: Breakdown by job role and department affiliation
Role Email Recipients Emails Opened Link Clicked Data Entered
Administration & Leadership 408 253 (62.01%) 45 (11.03%) 21 (5.15%)
Athletics 122 73 (59.84%) 12 (9.84%) 4 (3.28%)
Faculty & Research 1875 1272 (67.84%) 217 (11.57%) 88 (4.69%)
Finance 202 72 (35.64%) 26 (12.87%) 11 (5.45%)
Human Resources & Payroll 42 27 (64.29%) 6 (14.29%) 1 (2.38%)
IT & Information Security 201 107 (53.23%) 27 (13.43%) 10 (4.98%)
Student Roles 3193 1604 (50.23%) 449 (14.06%) 216 (6.76%)
Student Services & Advising 250 129 (51.60%) 28 (11.20%) 14 (5.60%)
Other Roles 2836 1054 (37.17%) 242 (8.53%) 98 (3.46%)
Total 9129 4591 (50.29%) 1052 (11.52%) 463 (5.07%)
TABLE VI: Effectiveness of phishing emails across different organizational roles. Percentages are relative to the number of email recipients.
To address RQ2, we analyzed the click-through rates across different job roles and department affiliations within the university. Table V presents the combined data for each department’s email activities, including the number of emails delivered, opened, links clicked, and data entered. The percentages provided are relative to the number of emails delivered to each department. Comparing the percentages across departments, we observe that the College of Business had the highest percentage of emails opened at 64.34%, followed by the College of Sciences at 60.47%. The College of Architecture, Construction and Planning had the highest percentage of links clicked at 13.25%, while the ”Other” category had the lowest at 8.49%. The Honors College and the College of Architecture, Construction and Planning had the highest percentages of data entered at 7.07% and 7.23%, respectively. Table VI shows the effectiveness of the phishing emails across different organizational roles. The percentages are relative to the number of email recipients for each role. Faculty & Research had the highest percentage of emails opened at 67.84%, while the ”Other Roles” category had the lowest at 37.17%. Human Resources & Payroll had the highest percentage of links clicked at 14.29%, followed closely by Student Roles at 14.06%. Student Roles, which include positions typically filled by students such as graduate assistants, teaching assistants, research assistants, and student workers, also had the highest percentage of data entered at 6.76%. The College of Business and the College of Sciences was observed to have higher email open rates, while the College of Architecture, Construction and Planning and the Honors College had higher percentages of links clicked and data entered. Similarly, Faculty & Research and Human Resources & Payroll roles exhibited higher engagement with phishing emails compared to other roles. The high engagement of Student Roles with phishing emails is particularly concerning, as these positions are often filled by less experienced individuals who may be more vulnerable to social engineering tactics.
VI-C RQ3: Motivation for engaging with phishing emails
To understand the underlying reasons and motivations for individuals to engage with LLM and human-written phishing emails, we conducted a questionnaire on 34 employees who entered data in the web interface after clicking the phishing link. This optional questionnaire was sent out to all 463 employees those who entered data, but only 34 completed the questionnaire. The questionnaire consisted of Likert scale questions and open-ended qualitative questions.
Likert Questions (Rating from 1-5: 1. Strongly Disagree, 2. Somewhat Disagree, 3. Neither Agree nor Disagree, 4. Somewhat Agree, 5. Strongly Agree):
  1. 1.
    ”The fact that the email came from an internal (@ORG_NAME.edu) email address influenced my decision to trust the email.” (Average likert score: 4.77)
  2. 2.
    ”I decided to interact with the email because I felt confident about the sender’s identity.” (Average likert score: 4.51)
  3. 3.
    ”My perception of the email’s relevance to my personal or professional interests significantly guided my decision to interact with it.” (Average likert score: 4.35)
  4. 4.
    ”My perception of the email’s urgency (such as a warning, a deadline, or a request) significantly guided my decision to interact with it.” (Average likert score: 3.43)
The high average likert scores for questions 1 and 2 suggest that the sender’s identity, particularly the fact that the email came from an internal email address, played a significant role in the respondents’ decision to trust and interact with the email. The high score for question 3 indicates that the perceived relevance of the email to the respondents’ personal or professional interests also influenced their decision to engage with the email. The lower score for question 4 suggests that while the perceived urgency of the email did influence the respondents’ decision, it was not as strong a factor as the sender’s identity and the email’s relevance.
Open-ended Qualitative Questions:
  1. 1.
    ”Which elements, if any, did you evaluate or scrutinize to assess the email’s legitimacy?”
    Only one respondent reported not scrutinizing the email, despite the fact that this survey was only taken by those who entered data into the phishing link. This could indicate either a lack of forthcomingness due to shame or other social stigma, or that the respondents proceeded to enter data despite scrutinizing the email’s legitimacy.
    Common elements scrutinized included the sender’s email address and the wording of the email. Some respondents reported looking for errors (typos, misspellings, or strange alphanumeric symbols) in the email body, as these are common in many phishing attacks.
  2. 2.
    ”What motivated you to engage with the email?”
    Respondents reported trusting the email since it came from an internal address and the link pointed to a domain name affiliated with the institution. This suggests an awareness of phishing but not of lateral phishing involving compromised internal accounts.
    The topic of the email was relevant to the recipient. For example, some respondents reported having discussed professional development with their supervisor earlier that day and then receiving the phishing template about professional development. Others found the timely email involving the solar eclipse convincing since they were aware of the institution’s planned activities related to the event.
    Some respondents reported being busy and stressed, and the urgency of the email prompted them to follow the instructions without further consideration.
    Some respondents felt pressure to complete the task in a timely manner since the email appeared to come from their supervisor.
  3. 3.
    ”Is there anything else you’d like to share about your experience with the email for us to better understand your interaction with the email?”
    Respondents mainly shared their thoughts and suggestions about the phishing training itself. Some appreciated the training and believed it would increase their phishing awareness in the future, while others expressed displeasure and considered it useless or a waste of time.
These findings provide valuable insights into the underlying reasons and motivations for individuals to engage with the lateral phishing emails. The respondents’ trust in internal email addresses, the perceived relevance of the email content, and the sense of urgency conveyed in the emails were key factors influencing their decision to interact with the phishing emails. The results also highlight the need for more comprehensive phishing awareness training that addresses lateral phishing and the exploitation of compromised internal accounts. Although some respondents were vigilant in identifying characteristics of typical phishing emails, they seemed vulnerable to the well-crafted emails that employed internal email addresses, suggesting a gap in their understanding of these more sophisticated phishing techniques that could be exacerbated in the age of LLM powered phishing. Additionally, the findings underscore the importance of fostering a culture of cybersecurity awareness and providing employees with the necessary tools and knowledge to identify and report suspicious emails, even when under pressure or stress.
VI-D Detection of LLM written phishing emails
Tagging LLM generated emails may be a solution to be explored in future work to help mitigate the threat of LLM phishing emails. When an LLM-generated email is detected, proactive measures may be taken to mitigate its potential impact. One potentially effective strategy is to automatically tag such emails, informing the recipient of the potential risks. This tag would act as an immediate visual cue, alerting the recipient that the email has been flagged as being potentially LLM-generated, thus warranting extra scrutiny. Email tagging has been investigated in previous research as it augments users ability to more quickly categorize content [50, 51]. Currently within the organization, a similar tagging methodology is used to indicate emails that come from outside the organization, assigning an [EXTERNAL] tag to all such emails. Such tags would not only serve as a direct warning but also reinforce the training employees receive, bridging the gap between theoretical knowledge and practical application. This dual approach of enhancing both technological defenses and employee vigilance could be part of a defense strategy against the evolving landscape of LLM-generated phishing threats. Future research could investigate email tagging for LLM-generated content to see if these tags could heighten user awareness to LLM phishing threats.
In order to investigate if LLM-generated phishing emails can be detected, we performed a small preliminary experiment using a recent LLM detection system that utilizes T5 [52] encoder embeddings in a Multilayer Perceptron (MLP) classifier [37]. We built our dataset by taking 5,000 human-written email samples [53] and generating corresponding machine-written variants using the Vicuna 13B v1.5 model, thus creating paired examples to capture the subtle differences between human and LLM-generated texts. The detection system works by feeding each email into a frozen T5 encoder that extracts dense, semantically rich embeddings, which are then processed by an MLP that learns to classify the texts; our preliminary results show that, after fine-tuning on this domain-specific dataset, the classifier achieves 98 F1 score on this test. It is important to note that this experiment is only a brief test within our broader study, and a more detailed investigation into advanced detection strategies remains a promising direction for future work.
While email tagging presents a promising defensive approach, several technical challenges must be addressed for effective implementation. First, detection mechanisms must balance sensitivity and specificity, where too aggressive detection could flag legitimate communications, and lenient systems might miss sophisticated LLM-generated phishing attempts. A fundamental technical challenge in the detection of LLM generated content lies in the detection of content from unknown or novel LLMs. While detecting generations from a specific, known LLM is relatively straightforward [35], real-world scenarios involve many different models (with tens of thousands of LLMs available on platforms like Hugging Face alone), making comprehensive coverage extremely difficult. This ”in-the-wild” detection problem (identifying LLM-generated content without prior knowledge of the specific generating model) remains largely unsolved and represents a significant barrier for detecting LLM generated text. False positives could erode user trust in the tagging system, potentially leading employees to ignore warnings altogether. Additionally, implementers must consider the computational overhead of real-time detection and the potential for attackers to adapt their strategies to evade detection. Future work should explore adversarial testing methodologies to continuously improve tagging systems against evolving LLM-based threats from both known and previously unseen models. Behavioral anomaly detection systems could monitor communication patterns, flagging deviations from established sender behaviors such as unusual timing, atypical requests, or shifts in writing style [54]. Organizations might also implement adaptive authentication protocols that escalate verification requirements when suspicious behaviors are detected [55]. These solutions should complement, rather than replace, user education focused specifically on identifying sophisticated LLM-generated phishing attempts. While email security solutions with advanced detection capabilities represent a significant initial investment, this cost must be contextualized against the potentially greater financial impact of security breaches. The long-term benefits of robust email security, including breach prevention and risk mitigation, can outweigh the upfront expenditure.
The threat landscape will likely evolve significantly as LLMs advance beyond current capabilities. Future iterations of models OpenAI’s GPT, Anthropic’s Claude, Google’s Gemini, and even open-source models such as Meta’s LLAMA models may be able to generate even more convincing phishing content with improved contextual understanding and reduced linguistic artifacts. Of particular concern is the emergence of multi-modal phishing attacks that combine text with deepfake audio or video elements [56]. For instance, a phishing email might include a fabricated video conference recording or voicemail from a supervisor, substantially increasing the attack’s credibility. These multi-modal approaches could exploit established trust relationships while circumventing traditional text-based detection methods.
VI-E Overall discussion and findings
@@
The after action review conducted by the cyber operations team yielded crucial insights into the employee preparedness in identifying phishing emails, particularly those generated by LLMs. Given the sophistication of LLM-generated emails, which often lack grammatical errors and appear highly credible, it’s imperative to enhance the security awareness training program. This updated training should focus on educating employees to discern subtle cues, such as the overall context of the email and the nature of the request. Alongside recognizing these nuances, it is equally important for employees to be alert to common phishing red flags like the urgency of requests and the origin of URLs or attachments. Table VI reveals that Student Roles were the most vulnerable job role in this lateral phishing exercise. This finding corroborates the observations made by the cyber operations team during real-world phishing attacks faced by the organization. The data suggests that focused phishing awareness training may be necessary for this more vulnerable population to effectively mitigate the risks associated with targeted phishing attempts.
Beyond identifying vulnerable populations, temporal factors may also significantly influence phishing susceptibility. Our conversations with the cyber operations team suggest that vulnerability likely fluctuates with organizational cycles and workload patterns. Specifically, the beginning of academic semesters represents a period of heightened vulnerability when many employees haven’t completed recent security training and when students are particularly susceptible to job-related phishing schemes. Other temporal risk factors include tax season, institutional financial deadlines, and periods following the publication of news releases featuring organizational members. These individuals often become immediate targets for personalized phishing attempts. While our current study design doesn’t capture these temporal variations, understanding these cyclical patterns could inform more strategic deployment of security resources and training refreshers during predicted high-risk periods.
Based on our findings, we propose several targeted enhancements to existing security awareness programs. Our data indicates that students workers are particularly vulnerable to phishing attempts. To address this vulnerability, we recommend a multi-faceted approach that leverages established social networks within the organization to disseminate educational content specifically tailored to at-risk populations. This approach would be complemented by a two-phase training structure consisting of comprehensive annual training followed by targeted refresher sessions that reinforce key concepts throughout the year. The establishment of student ambassador programs would further strengthen these initiatives by engaging representatives from the student population in both the ideation and feedback phases of security awareness development. The content of these enhanced training programs should emphasize that sophisticated phishing attempts may be professionally written and indistinguishable from legitimate communications from a stylistic perspective. This represents a significant shift from traditional training that often focuses on identifying grammatical errors or obvious inconsistencies. Scenario-based exercises simulating LLM-generated phishing attempts would train employees to assess contextual factors rather than superficial textual elements.
To measure the effectiveness of these enhanced training initiatives, we propose establishing a target phishing link click rate of less than 5% across the organization, with specific monitoring mechanisms for vulnerable populations such as students. Year-over-year comparative analysis using standardized phishing simulations (consistent topic, urgency and sender) would provide quantitative metrics of improvement, while qualitative feedback collected through student focus groups would offer insights into the efficacy of training and knowledge retention. While some of these trainings can largely be incorporated within existing training frameworks without significant additional expenditure, developing and delivering specialized training for specific subgroups may require additional resources. However, the targeted approach is justified by the disproportionate vulnerability demonstrated by certain populations and the potential organizational impact of successful phishing attacks involving these groups.
VI-F Limitations and Future Work
While our current study provides valuable insights into the effectiveness of LLM-generated phishing emails compared to human-crafted ones, we acknowledge several limitations that present opportunities for future research. First, our short-term experimental design could be expanded into longitudinal studies to assess the sustained effectiveness of countermeasures such as email tagging systems and specialized awareness training programs. Such long-term studies would help determine whether the initial awareness and vigilance developed through our proposed trainings persist over time or require periodic reinforcement. Second, as AI technologies evolve, the threat landscape extends beyond text-based phishing. Future work could examine multimodal phishing attacks that combine LLM-generated text with other AI-generated content such as deepfakes, synthetic voice (voice phishing or ”vishing”), or AI-generated images. These multi-vector attacks could potentially bypass traditional text-based detection methods and exploit different cognitive vulnerabilities in users. Third, our study measured credential entry as a binary outcome (entered vs. not entered), which provides clear effectiveness metrics but lacks granularity in understanding user hesitation behaviors. Future research could incorporate behavioral metrics such as session duration before submission, partial form completion patterns, and abandonment rates to better understand the decision-making process that occurs between receiving a phishing email and potential credential submission.
Our study’s focus on a higher education institution presents both advantages and limitations for generalizability to other types of organizations. Universities operate within unique organizational structures characterized by highly decentralized administration [61] similar to governmental organizations but distinct from many corporate environments, frequent personnel turnover particularly among student workers, and diverse demographic compositions across departments. Higher education institutions also face targeted phishing attempts due to the abundance of publicly available information about employees, making them attractive targets for social engineering [62]. While these characteristics provided a rich environment for our study, they also necessitate caution when extrapolating our findings to other organizational contexts. Future research could replicate our methodology across different sectors including corporate environments, healthcare institutions, and government agencies to establish the generalizability of our findings and develop context-specific defensive strategies that account for varying organizational structures, personnel dynamics, and security practices.
VII Conclusion
In this study, we have presented a comparative analysis of the effectiveness of lateral phishing emails generated by LLMs and those crafted by human communications professionals within a large educational organization. Our findings underscore the pressing threat posed by LLM-generated phishing emails, as they have proven to be as effective as, and in some cases even more effective than, their human-crafted counterparts. The experiments conducted across roughly 9,000 employees at the university revealed that these phishing emails can exploit trust relationships and power dynamics within the organization. Our analysis of click-through rates across different job roles and department affiliations identified potential patterns of vulnerability to lateral phishing attempts. These initial findings can help guide the development of awareness campaigns and the strategic allocation of resources to bolster defenses where they are needed most. The questionnaire administered to employees who entered data in the simulated phishing emails provided valuable insights into the underlying reasons and motivations for engaging with phishing content. The sender’s identity, the perceived relevance of the email, and the sense of urgency were identified as key factors influencing recipients’ decisions to interact with the emails. This highlights the need for comprehensive phishing awareness training that addresses lateral phishing and the exploitation of compromised internal accounts.
While the capability of LLMs to generate human-like text may seem intuitive, our study provides the first empirical quantification of this threat in a large-scale organizational context, offering insights for security practitioners on specific vulnerability patterns and countermeasures. This study contributes to the growing body of research on the emerging threat of LLM-generated phishing attacks and emphasizes the critical need for organizations to adapt their cybersecurity strategies accordingly. The findings underscore the importance of investing in robust employee training and awareness programs that specifically address the unique characteristics of AI-powered phishing attempts. Additionally, the insights gathered from this study can inform the development of advanced technological defenses, such as email tagging for LLM-generated content, to augment users’ ability to identify and avoid these threats. As the landscape of cybersecurity threats continues to evolve, organizations must remain vigilant and proactive in mitigating the risks posed by LLM-generated phishing attacks.